the following table gives the number of games won and lost by the winning team of a baseball league for eight consecutive years, How do i find the percentage won, for the 1st year won 53 lost 31, 2nd year won 57 lost 33, 3rd won 60 lost 32, 4th 55 won 41 lost, 5th 51 won 42 lost, 6th won 63 lost 39, 7th won 65 lost 40, and 8th year won 60 lost 45. how do i find the percentage won for each year.
Answer by Boreal(15235)

Percentage is win/total number of games
53/84=63.2%. 84/53.000... This is division. It goes .6
57/90=63.3%-------------------------------------504 subtract from 530 and get 26, bring down 0 and divide 84 into 260. It goes 3 times with 8 left over. Bring down the 0 and have 80, and 53 goes into it 1 time with 27 left over, which rounds to .2. All of these are done the same way. It is converting a fraction to a decimal.
60/92=65.2%
55/96=57.3%
51/93=54.8%
63/102=61.8%
65/105=61.9%
60/105=57.1%